# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v4

@@ I Introduction
To achieve Artificial General Intelligence (AGI), the development of embodied AI stands as a fundamental avenue. Different from conversational agents like ChatGPT [5], embodied AI believes that the true AGI can be achieved by controlling physical embodiments and interacting with both simulated and physical environments [6, 7, 8, 9, 10]. As we stand at the forefront of AGI-driven innovation, it is crucial to delve deeper into the realm of embodied AI, unraveling their complexities, evaluating their current developmental stage, and contemplating the potential trajectories they may follow in the future. Nowadays, embodied AI contains various key techniques across Computer Vision (CV), Natural Language Processing (NLP), and robotics, with the most representative being embodied perception, embodied interaction, embodied agents, and sim-to-real robotic control. Therefore, it is imperative to capture the evolving landscape of embodied AI in the pursuit of AGI through a comprehensive survey.
@@
The rest of this survey is organized as follows. Section 2 introduces various embodied robots. Section 3 describes general and real-scene embodied simulators. Section 4 introduces embodied perception, including active visual perception, 3D visual grounding, visual language navigation and non-visual perception. Section 5 introduces embodied interaction. Section 6 introduces embodied agents including the embodied multimodal foundation model and embodied task planning. Section 7 introduces sim-to-real adaptation including embodied world model, data collection and embodied control. Finally, we share several promising research directions in Section 8.
@@ II-A Fixed-base Robots
Fixed-base robots, as shown in Fig. 4 (a), are extensively employed in laboratory automation, educational training, and industrial manufacturing due to their compactness and high-precision operations. These robots feature robust bases and structures that ensure stability and high accuracy during operation. Equipped with high-precision sensors and actuators, they achieve micron-level precision, making them suitable for tasks that require high accuracy and repeatability [23]. Moreover, fixed-base robots are highly programmable, allowing users to adapt them for various task scenarios, such as Franka (Franka Emika panda) [24], Kuka iiwa (KUKA) [25], and Sawyer (Rethink Robotics) [26]. Despite their excellent performance in many fields, fixed-base robots have certain disadvantages. Firstly, their fixed-base design limits their operational range and flexibility, preventing them from moving or adjusting positions over large areas and leading to their collaboration with humans and other robots. Secondly, fixed-base robots are generally expensive and require specialized personnel for installation and maintenance, which increases their initial investment and operational costs [23, 27].
@@ II-B Wheeled Robots and Tracked Robots
In comparison, tracked robots, with powerful off-road capabilities and high maneuverability, show significant potential in agriculture, construction, and disaster recovery, as shown in Fig. 4 (c). The track system provides a larger ground contact area, distributing the robot’s weight and reducing the risk of sinking in soft terrain such as mud and sand. Moreover, tracked robots are typically equipped with robust power and suspension systems, allowing them to maintain stability and traction on complex terrains [30]. Consequently, reliable tracked robots are also used in sensitive areas such as the military. The iRobot PackBot is a versatile military-tracked robot capable of performing tasks such as reconnaissance, explosive ordnance disposal, and search and rescue missions [31]. However, due to the high friction of the track system, tracked robots often suffer from low energy efficiency. Additionally, their movement speed on flat surfaces is slower than wheeled robots, as well as the flexibility and maneuverability.
@@ II-C Quadruped Robots
Quadruped robots, known for their stability and adaptability, are well-suited for complex terrain exploration, rescue missions, and military applications. Inspired by quadrupedal animals, these robots can maintain balance and mobility on uneven surfaces, as shown in Fig. 4 (d). The multi-jointed design allows them to mimic biological movements, achieving complex gaits and posture adjustments. High adjustability enables the robots to automatically adapt their stance to changing terrain, enhancing maneuverability and stability. Sensing systems, such as LiDAR and cameras, provide environmental awareness, allowing the robots to navigate autonomously and avoid obstacles [32]. Researchers commonly use several types of quadruped robots as research platforms: Unitree Robotics, Boston Dynamics Spot, and ANYmal C. Unitree Robotics’ Unitree A1 and Go1 are noted for their cost-effectiveness and flexibility. The A1 [33] and Go1 [34] possess strong mobility and intelligent obstacle avoidance capabilities, suitable for various applications. Boston Dynamics’ Spot is renowned for its superior stability and operational flexibility, which are commonly used in industrial inspections and rescue missions. It features powerful load-carrying capacity and adaptability, capable of performing complex tasks in harsh environments [35]. ANYbotics’ ANYmal C, with its modular design and high durability, is widely employed in industrial inspection and maintenance. The ANYmal C is equipped with autonomous navigation and remote operation capabilities, suitable for prolonged outdoor tasks and even extreme lunar missions [36]. Like fixed-base robots, quadruped robots face similar drawbacks, such as high costs. The complex design and high manufacturing costs of quadruped robots result in substantial initial investments, limiting their use in cost-sensitive areas. Additionally, quadruped robots have limited battery endurance in complex environments, requiring frequent recharging or battery replacement for prolonged operation [37].
@@ II-D Humanoid Robots
Humanoid robots are distinguished by their human-like form and are increasingly prevalent in sectors such as the service industry, healthcare, and collaborative environments. These robots can mimic human movements and behavioral patterns, providing personalized services and support. Their dexterous hand designs enable them to perform intricate and complex tasks, distinguishing them from other types of robots, as shown in Fig. 4 (e). These hands typically have multiple degrees of freedom and high-precision sensors, allowing them to emulate the grasping and manipulation capabilities of human hands, which is particularly crucial in fields such as medical surgery and precision manufacturing [38, 39]. Among current humanoid robots, Atlas (Boston Dynamics) is renowned for its exceptional mobility and stability. Atlas can perform complex dynamic actions such as running, jumping, and rolling, demonstrating the potential of humanoid robots in highly dynamic environments [40]. The HRP series (AIST) is utilized in various research and industrial applications, with a design focus on high stability and flexibility, making it effective in complex environments, particularly for collaborative tasks with humans [41]. ASIMO (Honda), one of the most famous humanoid robots, can walk, run, climb stairs, and recognize faces and gestures, making it suitable for reception and guide services [42]. Additionally, a small social robot, Pepper (Softbank Robotics) can recognize emotions and engage in natural language communication and is widely used in customer service and educational settings [43].
@@ II-E Biomimetic Robots
Differently, biomimetic robots perform tasks in complex and dynamic environments by simulating efficient movements and functions of natural organisms. By emulating biological entities’ forms and movement mechanisms, these robots demonstrate significant potential in fields such as healthcare, environmental monitoring, and biological research [23]. Typically, they utilize flexible materials and structures to achieve lifelike, agile movements. These materials not only enhance the robots’ adaptability and flexibility but also minimize environmental impact. Furthermore, biomimetic robots are often equipped with advanced sensors and control systems, enabling real-time environmental sensing and rapid response, thereby enhancing their autonomous navigation and task execution capabilities. Importantly, biomimetic designs can significantly improve the robots’ energy efficiency by mimicking the efficient movement mechanisms of biological organisms, making them more economical regarding energy consumption [55, 56]. These biomimetic robots include fish-like robots [57, 58], insect-like robots [59, 60], and soft-bodied robots [61], as shown in Fig. 4 (f). However, biomimetic robots face several challenges. First, their design and manufacturing processes are complex and costly, limiting large-scale production and widespread application. Second, due to their use of flexible materials and complex movement mechanisms, the durability and reliability of biomimetic robots are limited in extreme environments.
@@ III-A General Simulator
The physical interactions and dynamic changes present in real environments are irreplaceable. However, deploying embodied models in the physical world often incurs high costs and faces numerous challenges. The ultimate goal of embodied AI is to transfer findings from virtual environments to real-world applications. Researchers can select simulators that best suit their needs to aid their studies. General-purpose simulators provide a virtual environment that closely mimics the physical world, allowing for algorithm development and model training, which offers significant cost, time, and safety advantages.
@@
Isaac Sim [45] developed by NVIDIA, is an advanced simulation platform tailored for robotics and AI research. The primary features of Isaac Sim include high-fidelity physical simulation, real-time ray tracing, an extensive library of robotic models, and deep learning support. Pixar’s USD (Universal Scene Description) format is also introduced to describe robots and complex scenes. Isaac Sim offers a variety of pre-built robotic models and environments, and it supports user-defined models. Its application scenarios include robotic navigation and control, autonomous driving, industrial automation, and human-robot interaction. By providing a robust and versatile platform, Isaac Sim significantly enhances the efficiency and effectiveness of robotics and AI research.
Gazebo [62] is an open-source simulator developed by Open Robotics, is widely used in robotics research and education. It offers high-fidelity physical simulation and rich features, making it a preferred tool for researchers and developers. Gazebo’s key features include high-fidelity physical simulation, diverse sensor simulation, extensive robot libraries, and tight integration with ROS. Gazebo supports the simulation of various sensors, including cameras, LiDAR, and sonar, and offers numerous pre-built robot models and environments with support for custom models. Its application scenarios include robot navigation and control and multi-robot systems.
PyBullet [48] is the Python interface for the Bullet physics engine, providing an easy-to-use simulation environment. PyBullet’s key features include ease of use, real-time physical simulation, diverse sensor simulation, and deep learning integration. PyBullet supports real-time physical simulation, including rigid body dynamics, collision detection, and constraint solving. Its application scenes include robot navigation and control, reinforcement learning, and computer graphics.
Table. II presents the key features and primary application scenarios of 10 general-purpose simulators. They each offer unique advantages in the field of embodied AI. Researchers can select the most appropriate simulator based on their specific research needs, thereby accelerating the development and application of embodied AI technologies. Fig. 5 shows the visualization effects of the general simulators.
@@ III-B Real-Scene Based Simulators
AI2-THOR [63] is an indoor embodied scene simulator based on Unity3D, launched in 2017 and led by the Allen Institute for Artificial Intelligence. As a high-fidelity simulator built in the real world, the biggest feature of AI2-THOR is its richly interactive scene objects and the physical properties assigned to them (such as open/close or even cold/hot). AI2-THOR consists of two parts: iTHOR and RoboTHOR. iTHOR contains 120 rooms categorized as kitchens, bedrooms, bathrooms, and living rooms, with over 2000 unique interactive objects, and supports multi-agent simulation; RoboTHOR contains 89 modular apartments with 600+ objects, the uniqueness of which is that these apartments correspond to real scenes in the real world. This means that researchers can remotely deploy their models in the real environment. So far, more than a hundred works have been published based on AI2-THOR.
@@
Matterport 3D [64] is proposed in R2R[65] in 2018, is more commonly used as a large-scale 2D-3D visual dataset rather than an embodied simulator. The Matterport3D dataset includes 90 architectural indoor scenes, comprises 10,800 panoramas and 194,400 RGB-D images, and provides surface reconstruction, camera posture, and 2D and 3D semantic segmentation annotations. In embodied AI, Matterport3D is mainly used for visual language navigation. Matterport3D transforms 3D scenes into discrete ‘viewpoints’, and embodied agents move between adjacent ‘viewpoints’ in Matterport3D scenes. At each ‘viewpoint’, embodied agents can obtain a 1280x1024 panorama image (18$\times$ RGB-D) centered on the ‘viewpoint’. The advantage of Matterport3D lies in its large, diverse, and detailed annotated 2D-3D data, as well as the ease of use and flexibility brought about by simulator-deploy-free. In the field of embodied navigation, Matterport3D is already one of the most important benchmarks.
@@
Virtualhome [66] is a home activity embodied AI simulator brought by Puig et al. in 2018. What makes Virtualhome special most is its environment represented by an environment graph. Specifically, the environment graph is a dictionary composed of nodes (corresponding to objects) and edges (corresponding to relationships). Nodes include all objects in the current environment, as well as corresponding ID, status, and other information. This kind of environment graph provides a new way for embodied agents to understand the environment. In addition, users can customize and modify the environment graph to achieve the custom configuration of scene objects. Similar to AI2-THOR, Virtualhome also provides a large number of interactive objects, and embodied agents can interact with them and change their status. Virtualhome also provides 6 humanoid agents that can be deployed at the same time and support the acquisition of RGB-D and semantic information from multiple perspectives. Another feature of Virtualhome is its simple and easy-to-use API. The actions of embodied agents are simplified to the format of ‘operation + object’. This feature makes Virtualhome widely used in the research fields of embodied planning, instruction decomposition, etc.
@@
Habitat [67] is an open-source simulator for large-scale human-robot interaction research launched by Meta in 2019. Habitat includes three parts: Habitat-sim, Habitat-lab, and Habitat-challenge. Among them, Habitat-sim is a high-performance 3D simulator based on the Bullet physics engine, which is the foundation of Habitat; Habitat-lab is an embodied simulation framework for reinforcement learning, encapsulated based on Habitat-sim; Habitat-challenge is a series of benchmarks based on Habitat. The biggest feature of Habitat is its extremely high degree of openness. Researchers can import and create their 3D scenes in Habitat or use the rich open resources on the Habitat platform for expansion. This gives Habitat several scenes that other simulation platforms cannot match. Habitat has a wealth of customizable sensors and supports multi-agent simulation. Multiple embodied agents from open resources or customizations (such as humans and robot dogs) can cooperate in the simulation scene, move freely, and perform simple interactions with the scene. Based on these advantages, Habitat is attracting increasing attention.
@@
Simulation Year Num of Continuous 3D Scene Sensors Physics Multiple Object 3D
Platform Scenes Action Scans Agents States Assets
@@
Different from other simulators that focus more on the scene, SAPIEN [68] pays more attention to simulating the interaction of objects in the scene. Based on the PhysX physics engine, SAPIEN provides fine-grained embodied control, which can implement joint control based on force and torque through the ROS interface, and the high-level action interface can also support the motion planning for embodied agents to avoid collisions. Based on the PartNet-Mobility Dataset, SAPIEN provides indoor simulation scenes containing rich interactive objects and supports the import of custom resources. Different from the API provided by simulators like AI2-THOR that directly changes the status of objects, SAPIEN supports simulated physical interactions, and embodied agents can control the hinged parts of objects through physical actions, thereby changing the status of objects. These features make SAPIEN very suitable for training the fine-grained object operation of embodied AI.
iGibson [69][70] is an open-source simulator launched by Stanford in 2021. iGibson is built on the Bullet physics engine, provides 15 high-quality indoor scenes, and supports the import of assets from other datasets such as Gibson and Matterport3D, enabling over 10,000 simulated scenes. As an object-oriented simulator, iGibson assigns rich changeable attributes to objects, not limited to the kinematic properties of objects (posture, speed, acceleration, joint configuration of articulated objects), but also includes temperature, humidity, cleanliness, switch status, etc. This allows iGibson to implement more complex embodied interactions and build more difficult, long-term embodied tasks. In addition, besides the depth and semantic sensors that are standard in other simulators, iGibson also provides LiDAR for embodied agents, allowing embodied agents to obtain 3D point clouds in the scene easily. Regarding embodied agent configuration, iGibson supports continuous action control and fine-grained joint control. This allows the embodied agents in iGibson to interact delicately with the objects in the scene while moving freely.
TDW [71] was launched by MIT in 2021. As one of the latest embodied simulators, TDW combines high-fidelity video and audio rendering, realistic physical effects, and a single flexible controller, making certain progress in the perception and interaction of the simulated environment. TDW integrates multiple physics engines into one framework, which can realize the physical interaction simulation of various materials such as rigid bodies, soft bodies, fabrics, and fluids, and provides situational sounds when interacting with objects. Thus, TDW has taken an important step compared to other simulators. TDW supports the deployment of multiple intelligent agents and provides users with a rich API library and asset library, allowing users to freely customize scenes and tasks according to their own needs, even outdoor scenes and related tasks. The extremely high degree of freedom gives TDW higher potential for embodied AI.

### v5

@@ I Introduction
To achieve Artificial General Intelligence (AGI), the development of embodied AI stands as a fundamental avenue. Different from conversational agents like ChatGPT [5], embodied AI believes that the true AGI can be achieved by controlling physical embodiments and interacting with both simulated and physical environments [6, 7, 8]. As we stand at the forefront of AGI-driven innovation, it is crucial to delve deeper into the realm of embodied AI, unraveling their complexities, evaluating their current developmental stage, and contemplating the potential trajectories they may follow in the future. Nowadays, embodied AI contains various key techniques across Computer Vision (CV), Natural Language Processing (NLP), and robotics, with the most representative being embodied perception, embodied interaction, embodied agents, and sim-to-real robotic control. Therefore, it is imperative to capture the evolving landscape of embodied AI in the pursuit of AGI through a comprehensive survey.
@@
The rest of this survey is organized as follows. Section 2 introduces various embodied robots. Section 3 describes general and real-scene embodied simulators. Section 4 introduces embodied perception, including active visual perception, 3D visual grounding, visual language navigation and non-visual perception. Section 5 introduces embodied interaction. Section 6 introduces embodied agents including the embodied multimodal foundation model and embodied task planning. Section 7 introduces sim-to-real adaptation including embodied world model, data collection and training, and embodied control. In Section 8, we discuss promising research directions.
@@ II-A Fixed-base Robots
Fixed-base robots, as shown in Fig. 4 (a), are extensively employed in laboratory automation, educational training, and industrial manufacturing due to their compactness and high-precision operations. These robots feature robust bases and structures that ensure stability and high accuracy during operation. Equipped with high-precision sensors and actuators, they achieve micron-level precision, making them suitable for tasks that require high accuracy and repeatability [21]. Moreover, fixed-base robots are highly programmable, allowing users to adapt them for various task scenarios, such as Franka (Franka Emika panda) [22], Kuka iiwa (KUKA) [23], and Sawyer (Rethink Robotics) [24]. Nevertheless, fixed-base robots have certain disadvantages. Their fixed-base design limits their operational range and flexibility, preventing them from moving or adjusting positions over large areas and leading to their collaboration with humans and other robots.  [21].
@@ II-B Wheeled Robots and Tracked Robots
Differently, tracked robots have powerful off-road capabilities and maneuverability, showing potential in agriculture, construction, and disaster recovery, as shown in Fig. 4 (c). The track system provides a larger ground contact area, distributing the robot’s weight and reducing the risk of sinking in soft terrain such as mud and sand. Moreover, tracked robots are equipped with robust power and suspension systems, to maintain stability and traction on complex terrains [27]. Consequently, tracked robots are also used in sensitive areas such as the military. The iRobot’s PackBot is a versatile military-tracked robot capable of performing tasks such as reconnaissance, explosive ordnance disposal, and rescue missions [28]. However, due to the high friction of the track system, tracked robots often suffer from low energy efficiency. Additionally, their movement speed on flat surfaces is slower than wheeled robots, as well as their flexibility and maneuverability.
@@ II-C Quadruped Robots
Quadruped robots, known for their stability and adaptability, are well-suited for complex terrain exploration, rescue missions, and military applications. Inspired by quadrupedal animals, these robots can maintain balance and mobility on uneven surfaces, as shown in Fig. 4 (d). The multi-jointed design allows them to mimic biological movements, achieving complex gaits and posture adjustments. High adjustability enables the robots to automatically adapt their stance to changing terrain, enhancing maneuverability and stability. Sensing systems, such as LiDAR and cameras, provide environmental awareness, allowing the robots to navigate autonomously and avoid obstacles [29]. Several types of quadruped robots are widely used: Unitree Robotics, Boston Dynamics Spot, and ANYmal C. Unitree Robotics’ Unitree A1 and Go1 are noted for their cost-effectiveness and flexibility. The A1 [30] and Go1 [31] possess strong mobility and intelligent obstacle avoidance capabilities, suitable for various applications. Boston Dynamics’ Spot is renowned for its superior stability and operational flexibility, which are commonly used in industrial inspections and rescue missions. It features powerful load-carrying capacity and adaptability, capable of performing complex tasks in harsh environments [32]. ANYbotics’ ANYmal C, with its modular design and high durability, is widely employed in industrial inspection and maintenance. The ANYmal C is equipped with autonomous navigation and remote operation capabilities, suitable for prolonged outdoor tasks and even extreme lunar missions [33]. The complex design and high manufacturing costs of quadruped robots result in substantial initial investments, limiting their use in cost-sensitive areas. Additionally, they have limited battery endurance in complex environments, requiring frequent recharging or battery replacement for prolonged operation [34].
@@ II-D Humanoid Robots
Humanoid robots are distinguished by their human-like form and are increasingly prevalent in sectors such as the service industry, healthcare, and collaborative environments. These robots can mimic human movements and behavioral patterns, providing personalized services and support. Their dexterous hand designs enable them to perform intricate and complex tasks, distinguishing them from other types of robots, as shown in Fig. 4 (e). These hands typically have multiple degrees of freedom and high-precision sensors, allowing them to emulate the grasping and manipulation capabilities of human hands, which is particularly crucial in fields such as medical surgery and precision manufacturing [35]. Among current humanoid robots, Atlas (Boston Dynamics) is renowned for its exceptional mobility and stability. Atlas can perform complex dynamic actions such as running, jumping, and rolling, demonstrating the potential of humanoid robots in highly dynamic environments [36]. The HRP series (AIST) is utilized in various research and industrial applications, with a design focus on high stability and flexibility, making it effective in complex environments, particularly for collaborative tasks with humans [37]. ASIMO (Honda), one of the most famous humanoid robots, can walk, run, climb stairs, and recognize faces and gestures, making it suitable for reception and guide services [38]. Additionally, a small social robot, Pepper (Softbank Robotics) can recognize emotions and engage in natural language communication and is widely used in customer service and educational settings [39].
@@ II-E Biomimetic Robots
Differently, biomimetic robots perform tasks in complex and dynamic environments by simulating efficient movements and functions of natural organisms. By emulating biological entities’ forms and movement mechanisms, these robots demonstrate significant potential in fields such as healthcare, environmental monitoring, and biological research [21]. Typically, they utilize flexible materials and structures to achieve lifelike, agile movements and minimize environmental impact. Importantly, biomimetic designs can significantly improve the robots’ energy efficiency by mimicking the efficient movement mechanisms of biological organisms, making them more economical regarding energy consumption [42, 43]. These biomimetic robots include fish-like robots [44, 45], insect-like robots [46, 47], and soft-bodied robots [48], as shown in Fig. 4 (f). However, biomimetic robots face several challenges. First, their design and manufacturing processes are complex and costly, limiting large-scale production and widespread application. Second, due to their use of flexible materials and complex movement mechanisms, the durability and reliability of biomimetic robots are limited in extreme environments.
@@ III Embodied Simulators
TABLE II: General Simulator. HFPS: high-fidelity physical simulation; HQGR: high-quality graphics rendering; RRL: rich robot library; DLS: deep learning support; LSPC: large-scale parallel computing; ROS: tight integration with ROS; MSS: multiple sensor simulation; CP: cross-platform Nav: robot navigation AD: auto driving; RL: reinforcement learning LSPS: large-scale parallel sim MR: multi-robot systems RS: robot simulation. $\circ$ indicates that the simulator excels at this aspect.
Isaac Sim [49] 2023 $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ PhysX Nav, AD
Isaac Gym [50] 2019 $\circ$ $\circ$ $\circ$ PhysX RL,LSPS
Gazebo [51] 2004 $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ ODE, Bullet, Simbody, DART Nav,MR
PyBullet [52] 2017 $\circ$ $\circ$ Bullet RL,RS
Webots [53] 1996 $\circ$ $\circ$ $\circ$ $\circ$ ODE RS
MuJoCo [54] 2012 $\circ$ $\circ$ $\circ$ Custom RL, RS
Unity ML-Agents [55] 2017 $\circ$ $\circ$ $\circ$ Custom RL, RS
AirSim [56] 2017 $\circ$ $\circ$ Custom Drone sim, AD, RL
MORSE [57] 2015 $\circ$ $\circ$ Bullet Nav, MR
CoppeliaSim (V-REP) [58] 2013 $\circ$ $\circ$ $\circ$ $\circ$ Bullet, ODE, Vortex, Newton MR, RS
@@ III-A General Simulator
The physical interactions and dynamic changes present in real environments are irreplaceable. However, deploying embodied models in the physical world often incurs high costs and faces numerous challenges. General-purpose simulators provide a virtual environment that closely mimics the physical world, allowing for algorithm development and model training, which offers significant cost, time, and safety advantages.
@@
Isaac Sim [49] is an advanced simulation platform for robotics and AI research. It has high-fidelity physical simulation, real-time ray tracing, an extensive library of robotic models, and deep learning support. Its application scenarios include autonomous driving, industrial automation, and human-robot interaction.Gazebo [59] is an open-source simulator for robotics research. It has extensive robot libraries, and tight integration with ROS. It supports the simulation of various sensors and offers numerous pre-built robot models and environments. It is mainly used for robot navigation and control and multi-robot systems.PyBullet [52] is the python interface for the Bullet physics engine. It is easy to be used and has diverse sensor simulation and deep learning integration. PyBullet supports real-time physical simulation, including rigid body dynamics, collision detection, and constraint solving. Table. II presents the key features and primary application scenarios of 10 general-purpose simulators. They each offer unique advantages in the field of embodied AI. Researchers can select the most appropriate simulator based on their specific research needs, thereby accelerating the development and application of embodied AI technologies. Fig. 5 shows the visualization effects of the general simulators.
@@ III-B Real-Scene Based Simulators
AI2-THOR [60] is an indoor embodied scene simulator based on Unity3D, led by the Allen Institute for Artificial Intelligence. As a high-fidelity simulator built in the real world, AI2-THOR has richly interactive scene objects and the physical properties assigned to them (such as open/close or even cold/hot). AI2-THOR consists of two parts: iTHOR and RoboTHOR. iTHOR contains 120 rooms categorized as kitchens, bedrooms, bathrooms, and living rooms, with over 2000 unique interactive objects, and supports multi-agent simulation; RoboTHOR contains 89 modular apartments with 600+ objects, the uniqueness of which is that these apartments correspond to real scenes in the real world. So far, more than a hundred works have been published based on AI2-THOR.
@@
Matterport 3D [61] is proposed in R2R[62] and is more commonly used as a large-scale 2D-3D visual dataset. The Matterport3D dataset includes 90 architectural indoor scenes, comprises 10,800 panoramas and 194,400 RGB-D images, and provides surface reconstruction, camera posture, and 2D and 3D semantic segmentation annotations. Matterport3D transforms 3D scenes into discrete “viewpoints”, and embodied agents move between adjacent “viewpoints” in Matterport3D scenes. At each “viewpoint”, embodied agents can obtain a 1280x1024 panorama image (18$\times$ RGB-D) centered on the “viewpoint”. Matterport3D is one of the most important embodied navigation benchmarks.
@@
Virtualhome [63] is a home activity embodied AI simulator brought by Puig et al. What makes Virtualhome special most is its environment represented by an environment graph. The environment graph represents the objects in the scene and their related relationships. Users can also customize and modify the environment graph to achieve the custom configuration of scene objects. This kind of environment graph provides a new way for embodied agents to understand the environment. Similar to AI2-THOR, Virtualhome also provides a large number of interactive objects, and embodied agents can interact with them and change their status. Another feature of Virtualhome is its simple and easy-to-use API. The actions of embodied agents are simplified to the format of “operation + object”. This feature makes Virtualhome widely used in the research fields of embodied planning, instruction decomposition, etc.
@@
Habitat [64] is an open-source simulator for large-scale human-robot interaction launched by Meta. Based on the Bullet physics engine, Habitat has implemented high-performance, high-speed, parallel 3D simulation, and provides a rich interface for reinforcement learning of embodied agents. Habitat has extremely high degree of openness. Researchers can import and create 3D scenes in Habitat or use the rich open resources on the Habitat platform for expansion. Habitat has many customizable sensors and supports multi-agent simulation. Multiple embodied agents from open resources or customizations (e.g., humans and robot dogs) can cooperate in the simulator, move freely, and perform simple interactions with the scene. Thus, Habitat is attracting increasing attention.
@@
Simulation Platform Year Num of Scenes Continuous Action 3D Scene Scans Sensors Physics Multiple Agents Object States 3D Assets
@@
Different from other simulators that focus more on the scene, SAPIEN [65] pays more attention to simulate the interaction of objects. Based on the PhysX physics engine, SAPIEN provides fine-grained embodied control, which can implement joint control based on force and torque through the ROS interface. Based on the PartNet-Mobility Dataset, SAPIEN provides indoor simulation scenes containing rich interactive objects and supports the import of custom resources. Different from simulators like AI2-THOR that directly changes the status of objects, SAPIEN supports simulated physical interactions, and embodied agents can control the hinged parts of objects through physical actions, thereby changing the status of objects. These features make SAPIEN very suitable for training the fine-grained object operation of embodied AI.
iGibson [66][67] is an open-source simulator launched by Stanford. Built on the Bullet physics engine, iGibson provides 15 high-quality indoor scenes, and supports the import of assets from other datasets such as Gibson and Matterport3D. As an object-oriented simulator, iGibson assigns rich changeable attributes to objects, not limited to the kinematic properties of objects (posture, speed, acceleration, etc.), but also includes temperature, humidity, cleanliness, switch status, etc. In addition, besides the depth and semantic sensors that are standard in other simulators, iGibson also provides LiDAR for embodied agents, allowing agents to obtain 3D point clouds in the scene easily. Regarding embodied agent configuration, iGibson supports continuous action control and fine-grained joint control. This allows the embodied agents in iGibson to interact delicately with the objects while moving freely.
TDW [68] was launched by MIT. As one of the latest embodied simulators, TDW combines high-fidelity video and audio rendering, realistic physical effects, and a single flexible controller, making certain progress in the perception and interaction of the simulated environment. TDW integrates multiple physics engines into one framework, which can realize the physical interaction simulation of various materials such as rigid bodies, soft bodies, fabrics, and fluids, and provides situational sounds when interacting with objects. Thus, TDW has taken an important step compared to other simulators. TDW supports the deployment of multiple intelligent agents and provides users with a rich API library and asset library, allowing users to freely customize scenes and tasks according to their own needs, even outdoor scenes and related tasks.